the perimeter of a rectangle is 94m. The length is 8m more than twice the width. Find the dimensions
Answer by mananth(16949)

Perimeter = 94
Length 1 ( 2 x + 8 )
Width = x ft

Perimeter =2((2x +8)+x)
Perimeter =2(2x+8+x)
Perimeter =4x+16 +2x
94=6x+16

94-16=6x
78=6x
x=13 m Width
Length =34 m